Consider a firm with the following production function q = 10 l1/2 k1/2. It aims to produce 1000 units of output and faces prices for labour and capital as follows: w =20, v =5


1
Expert's answer
2021-01-11T07:48:22-0500

Having this kind of question, we can do the following;

a)      Form the Lagrangian Function for the the cost minimization problem.  

solution

production function, "q=10 l^{1\/2} k^{1\/2}"

total output "=1000units"

price of labor, "w=20"

price of capital, "v=5"

total cost"=wl+vk=20l+5k"

Minimizing cost subject to the production function by setting up Lagrange:

a) min "wl+vk" "st" "q"

"L=wl+vk+\lambda(q-10l^{1/2}k^{1/2}"

b)      Find the cost-minimizing bundle of labor and capital, (k*, l*). 

solution

Solving Lagrange: Differentiating and equating it to zero

"dL/dl=w-\lambda 5l^{-1/2}k^{1/2}=0"

"w=\lambda5l^{-1/2}k^{1/2}...............(1)"

"dL/dk=v-\lambda5k^{-1/2}l^{1/2}=0"

"v=\lambda5k^{-1/2}l^{1/2}............(2)"

"dL/d\lambda=q-10l^{1/2}k^{1/2}=0"

"q=10l^{1\/2}k^{1\/2}............(3)"

Dividing (1) and(2)

"w\/v=k\/l"

"k=l(w\/v)"

Putting this in (3)

"q=10l^{1\/2}(l(w\/v))^{1\/2}=10l(w\/v)^{1\/2}"

"l^{*}=[q(v\/w)^{1\/2}]\/10"

Putting l in equation of k

"k=l(w\/v)=(w\/v)[q(v\/w)^{1\/2}]\/10"

"k^{*}=[q(v\/w)^{1\/2}]\/10"